OpenStudy (anonymous):

Find the equation in standard form which passes through (3, –6) and (–2, –1). You must show all work.

OpenStudy (anonymous):

@apoorvk may you help me with this ?

OpenStudy (apoorvk):

For a line passing through two points say, \((x_1,y_1)\) and \((x_2,y_2)\), the equation is: \[\frac{y-y_1}{x-x_1} = \frac{y_1 - y_2}{x_1 - x_2} \] Can you use this for the points given to you? (after this, we will rearrange it in standard form)

OpenStudy (anonymous):

I think the second equation will be -1

OpenStudy (apoorvk):

you mean the RHS? yes that would be '-1'. wbu, the LHS?

OpenStudy (apoorvk):

*what about the LHS